five years ago Tom was one third as old as his father was one third as old as his father wa then. In 5 years Tom will be half as old as his father will be then. Find their ages now.
Answer by stanbon(75887)

five years ago Tom was one third as old as his father was then.
In 5 years Tom will be half as old as his father will be then.
Find their ages now
------------
Equations:
t-5 = (1/3)(f-5)
t+5 = (1/2)(f+5)
---------
Modify::
3t-15 = f-5
2t+10 = f+5
------
Subtract and solve for "t"::
t-25 = -10
Tom = 15 yrs (Tom's age now)
----
Solve for "f"::
f+5 = 2t+10
f+5 = 30+10
Father = 35 yrs (Father's age now)
----------
Cheers,
Stan H.
------------------